Looking for values of all.
interest Caklate the value of each ofthe bonds shoan nhe rgubl·alof.hch pay nterest senanvaly qCck onthe Par Value 12 12 The value of bond A is(Round to th 1 2 3 578 9 0
Answer #1

A.Given:

Future Value (FV)= $500

Coupon rate= 9%/2= 4.5%

Coupon payment= 0.045*500= 22.50

Time period= 11 years*2= 22 semi-annual periods

Interest rate= 10%/2= 5%

The value of the bond is calculated by computing the present value of the bond.

The value of the bond is calculated using a financial calculator by inputting the below:

FV= 500; PMT= 22.50; N= 22; I/Y= 5

Press CPT and PV to calculate the present value of the bond.

The present value of bond A is $467.09

B.Given:

Future Value (FV)= $500

Coupon rate= 12%/2= 6%

Coupon payment= 0.06*500= 30

Time period= 15 years*2= 30 semi-annual periods

Interest rate= 12%/2= 6%

The value of the bond is calculated using a financial calculator by inputting the below:

FV= 500; PMT= 30; N= 30; I/Y= 6

Press CPT and PV to calculate the present value of the bond.

The present value of bond B is $500.

C.Given:

Future Value (FV)= $100

Coupon rate= 16%/2= 8%

Coupon payment= 0.16*100= 16

Time period= 4 years*2= 8 semi-annual periods

Interest rate= 16%/2= 8%

The value of the bond is calculated using a financial calculator by inputting the below:

FV= 100; PMT= 16; N= 8; I/Y= 8

Press CPT and PV to calculate the present value of the bond.

The present value of bond C is $145.97.

I hope that was useful :)